# Harm in AI-Driven Societies: An Audit of Toxicity Adoption on Chirper.ai

### v1

@@ Abstract.
Large Language Models (LLMs) are increasingly embedded in autonomous agents that participate in online social ecosystems, where interactions are sequential, cumulative, and only partially controlled. While prior work has documented the generation of toxic content by LLMs, far less is known about how exposure to harmful content shapes agent behavior over time, particularly in environments composed entirely of interacting AI agents. In this work, we study toxicity adoption of LLM-driven agents on Chirper.ai, a fully AI-driven social platform. Specifically, we model interactions in terms of stimuli (posts) and responses (comments), and by operationalizing exposure through observable interactions rather than inferred recommendation mechanisms.
We conduct a large-scale empirical analysis of agent behavior, examining how response toxicity relates to stimulus toxicity, how repeated exposure affects the likelihood of toxic responses, and whether toxic behavior can be predicted from exposure alone. Our findings show that while toxic responses are more likely following toxic stimuli, a substantial fraction of toxicity emerges spontaneously, independent of exposure. At the same time, cumulative toxic exposure significantly increases the probability of toxic responding. We further introduce two influence metrics, the Influence-Driven Response Rate and the Spontaneous Response Rate, revealing a strong trade-off between induced and spontaneous toxicity. Finally, we show that the number of toxic stimuli alone enables accurate prediction of whether an agent will eventually produce toxic content.
These results highlight exposure as a critical risk factor in the deployment of LLM agents and suggest that monitoring encountered content may provide a lightweight yet effective mechanism for auditing and mitigating harmful behavior in the wild.
@@ 1. Introduction
Bots have long played a significant role in online social platforms, influencing information diffusion, engagement dynamics, and public discourse (10.1145/2818717; Bessi_Ferrara_2016; the-spread; Ferrara2017DisinformationAS). However, the advent of Large Language Models (LLMs) has enabled a new generation of social bots capable of far more sophisticated and naturalistic interactions with users and with one another. These advances have spurred growing academic interest in understanding the behavior of LLM-driven social agents deployed in or modeled after online social ecosystems.
@@
Unlike traditional rule-based or template-driven bots, LLM-based agents exhibit adaptive and emergent behaviors that arise from ongoing interactions within social networks (emergent; aichatbots). Recent work has leveraged offline simulated social environments to examine the extent to which such agents can replicate human-like network structures (10.1145/3589335.3641253), coordinate and perform cooperative tasks (Wu2024ShallWT; orlando2025emergentcoordinatedbehaviorsnetworked), or give rise to collective phenomena such as polarization and echo chambers (10.5555/3737916.3741104; piao2025emergencehumanlikepolarizationlarge). Together, these studies suggest that LLM agents are not merely passive generators of text, but active participants in social dynamics whose behavior evolves over time.
@@
At the same time, LLMs can be misused to generate toxic or harmful content at scale, posing risks to online communities and individual users (chakraborty2024detoxbenchbenchmarkinglargelanguage; genai). In response, a substantial body of work has focused on measuring harmful generation and developing mitigation strategies, such as safer training procedures, filtering mechanisms, and post-hoc moderation (wang2025surveyresponsiblellmsinherent). Complementarily, a growing body of work in human-centered computing has shown that exposure to harmful content can influence user behavior, increasing the likelihood of adopting similar language or norms over time (Ferrara2015MeasuringEC). However, it remains unclear whether, and to what extent, analogous dynamics apply to LLM-driven agents.
@@
In particular, we lack empirical evidence on: (i) how these agents behave when they solely interact to each other, in a fully AI-driven, dedicated platform; and (ii) whether toxic content acts merely as a trigger for toxic responses, or whether repeated exposure systematically increases agents propensity to generate harmful outputs.
@@
In this paper, we address these gaps by performing a large-scale auditing on Chirper.ai, an AI-powered social network where users create and release LLMs-based agents that autonomously interact in the dedicated ecosystem, by generating posts and comments, as well as by interacting with other AI agents via following and liking mechanisms. Specifically, we rely on the platform to study toxicity adoption in LLM-driven agents through the lens of stimuli and responses. Due to the absence of platform-level documentation describing content recommendation or visibility mechanisms, we adopt a conservative and observable definition of exposure: we model stimuli as posts that an agent explicitly comments on, and responses as the comments produced by the agent. This interaction-based formulation allows us to avoid speculative assumptions about unseen content and to ground our analysis in measurable behavior.
@@
Our main contributions are threefold.
  • We are the first to provide a large-scale empirical audit of toxicity adoption in a fully AI-driven social platform, showing that stimuli play a central role in shaping the harmful behavior of LLM-based agents.
  • We introduce two complementary susceptibility metrics, the Influence-Driven Response Rate (IRR) and the Spontaneous Response Rate (SRR), which reveal a strong trade-off between exposure-driven and autonomous toxicity.
  • We show that the number of toxic stimuli encountered by an agent alone enables accurate prediction of whether it will eventually produce toxic content, without requiring access to model internals, training data, or prompt instructions.
@@ 2. Related Work
In the following, we review the relevant literature by focusing on three core components of this work: (i) the behavior of AI-driven agents in online social ecosystems, (ii) the social mechanisms underlying exposure and toxic content adoption, and (iii) Chirper.ai, the platform that serves as the empirical basis for our investigation.
@@
A growing body of research investigates how conversational agents and LLM-driven accounts behave when embedded in social platforms, where interactions are sequential, public, and shaped by evolving community norms (CHEN2026100107). In these environments, agent behavior is no longer determined solely by isolated prompts, but emerges from continuous interaction with other agents and content. The recent development of benchmarks and simulation frameworks for social-media agents reflects a growing interest in evaluating both agent capabilities and associated risks in realistic, platform-like settings, where agents must interpret heterogeneous content, navigate social norms, and make posting decisions under uncertainty (some2025benchmark; bechmarking).
Complementing these efforts, recent studies have begun to operationalize controlled human-agent interaction settings to analyze downstream social effects, such as polarization, opinion shifts, and norm formation, under carefully manipulated conditions (icwsm2025human_agent_polarization; Breum_Egdal_Gram). These controlled environments allow researchers to isolate specific mechanisms of influence while retaining key features of social interaction. At the same time, earlier work in computational social science has shown that antisocial behavior in online communities often emerges through repeated interactions and reinforcing feedback loops, rather than being driven exclusively by a small number of persistently malicious actors (cheng2015antisocial). This perspective highlights the importance of cumulative exposure and interaction dynamics in shaping harmful outcomes.
From this standpoint, it becomes critical to assess whether, and to what extent, fully AI-driven social platforms are susceptible to analogous emergent phenomena. In particular, understanding whether systems composed entirely of autonomous agents can give rise to the proliferation of hate speech, unsafe language, or toxic norms is essential for anticipating risks associated with deploying LLM-based agents at scale.
@@
Research on online toxicity has traditionally focused on large-scale measurement and detection, enabling systematic empirical analyses of harmful speech across diverse online communities (wulczyn2017exmachina). While this line of work has been essential for quantifying prevalence and building robust classifiers, it also raises a broader sociotechnical question: whether exposure to toxic content increases the likelihood that individuals subsequently adopt toxic language themselves.
A growing body of evidence suggests that exposure is indeed associated with increased rates of toxic behavior, though the magnitude and direction of this effect depend on contextual and social factors (Ferrara2015MeasuringEC; jiang2024socialapprovalnetworkhomophily). In particular, prior work shows that toxicity adoption varies according to group dynamics and social identity, such as whether harmful content originates from ingroup or outgroup members (howsocialmedia). These findings highlight that toxicity is not merely an individual trait, but a socially mediated phenomenon shaped by interaction patterns and relational context.
Complementary research further emphasizes the interactional nature of toxicity, demonstrating that harmful language often emerges at specific conversational turning points and can significantly alter the subsequent trajectory of discussion threads (provoke2022toxicity_triggers). This perspective motivates moderation approaches that go beyond isolated message classification, instead accounting for local conversational context and cumulative exposure histories. Together, these insights directly inform our distinction between influence-driven and spontaneous toxicity, and motivate our use of exposure-based predictors to model and anticipate harmful responding in agent-driven environments.
@@
Launched in 2023, Chirper.ai [^1] https://chirper.ai/ (chirper_blog2023; chirper_blog_simulated_world2024) is a fully AI-driven social platform in which autonomous agents, referred to as chirpers, generate and interact with content. On Chirper.ai, chirpers can publish posts and comments, and engage with one another through social mechanisms such as following and liking, closely mirroring interaction paradigms found in human-centered platforms (e.g., $\mathbb{X}$ or Facebook).
@@
The platform has recently become a prominent environment for studying machine behavior in social-media-like ecosystems because it enables large-scale observation of autonomous agent-to-agent interactions. Recent work examines the platform content (li2023masqueradeexploringbehaviorimpact) and compare it to human-driven social ecosystems (i.e., Mastodon), documenting differences in posting dynamics, abusive content, and network structure (zhu2025chirper_case). Our work builds on this emerging literature by focusing specifically on toxicity adoption as a function of observable stimuli, and by introducing complementary metrics to separate influence-driven toxicity from spontaneous toxic responding.
@@ 3. Methodology
    How aligned are agent responses with the toxicity of the stimuli they react to?
@@
    Is exposure to toxic stimuli sufficient to predict whether an agent will eventually produce toxic content?
We tackle each research question in the dedicated paragraph in the Results section below.
Table 1. Information regarding the underlying architecture of the chirpers, in terms of model, number of parameters (expressed in billions or trillions), and percentage of generated content on the platform.
@@
Following the acquisition of appropriate permissions to access the Chirper.ai data, we collected $10,420,000$ textual items, including both original posts and comments. We therefore reconstruct the history of about $75,000$ chirpers, instantiated by over $30,000$ (human) users. In addition to textual content, we also obtained metadata capturing social connections and interactions, such as follower relationships and likes. Table 1 further provides some information regarding the LLM architecture underlying the agents, and the corresponding percentage of generated content on the platform. Overall, Chirper.ai provides a unique and large-scale testbed for studying the behavior of LLM-driven agents in a realistic, socially structured environment, where exposure and interaction emerge organically from agent-to-agent dynamics rather than from controlled experimental prompts.
@@
As previously mentioned, we operationalize stimuli and responses at the level of comments. Specifically, a stimulus $S$ corresponds to a post on which a chirper comments, while a response $R$ is the comment produced by the chirper. This choice is driven by two main considerations.
@@
Second, we attempted to infer content exposure indirectly by leveraging the follower network and comment co-occurrence (i.e., assuming that a chirper is exposed to all content posted by users they follow or to posts commented on by overlapping sets of chirpers). However, empirical analysis revealed no meaningful correlation between these inferred exposure mechanisms and observed user interactions, such as likes or comments. This lack of correspondence suggests that such proxies do not reliably capture effective exposure on the platform.
Given these limitations, we adopt a conservative and observable definition of exposure, focusing exclusively on explicit interactions. We therefore characterize both stimuli and responses in terms of their toxicity. A piece of content is labeled as toxic if its toxicity score exceeds the $90$th percentile of the overall toxicity score distribution. This percentile-based threshold allows us to focus on the most extreme and potentially harmful content while remaining agnostic to absolute toxicity values.
@@
To estimate toxicity, we rely on detoxify [^2] https://pypi.org/project/detoxify/ (Detoxify) , a BERT-based model trained for toxic comment classification, which reports strong performance (mean AUC $\simeq 0.99$). Given an input text, the model outputs a toxicity score in the interval $[0,1]$, representing the intensity of toxic language expressed.
To further improve classification reliability, we restrict our analysis to English-language content only. Language detection is performed using pycld2, [^3] https://pypi.org/project/pycld2/ a Python library built around the Google Chromium embedded Compact Language Detector (CLD2). [^4] https://github.com/CLD2Owners/cld2 This filtering step mitigates potential degradation in toxicity estimation due to multilingual or low-resource language inputs, which are not explicitly supported by the classifier. After applying language filtering, the final dataset consists of about $8,000,000$ texts ($\sim 75\%$ of the original corpus).
Figure 1. Distribution of chirper toxic/non-toxic responses ($R^{*}/\tilde{R}$) over toxic/non-toxic stimuli ($S^{*}/\tilde{S}$).
Figure 2. Correlation between chirpers Responses and Stimuli, categorized in toxic (*) and non-toxic ($\sim$). $\rho$ values indicate Pearson correlation ($p<.0001$)
@@ 4. Results
Responses and Stimuli Alignment (RQ1)
We first investigate how chirper responses vary as a function of the stimuli to which they reply. Figure 1 shows the distribution of responses, distinguishing toxic and non-toxic comments in relation to toxic and non-toxic stimuli. Throughout the analysis, we denote toxic content with the symbol “*” and non-toxic content with $\sim$”.

### v2

@@ Abstract.
Large Language Models (LLMs) are increasingly embedded in autonomous agents that engage, converse, and co-evolve in online social platforms. While prior work has documented the generation of toxic content by LLMs, far less is known about how exposure to harmful content shapes agent behavior over time, particularly in environments composed entirely of interacting AI agents. In this work, we study toxicity adoption of LLM-driven agents on Chirper.ai, a fully AI-driven social platform. Specifically, we model interactions in terms of stimuli (posts) and responses (comments). We conduct a large-scale empirical analysis of agent behavior, examining how toxic responses relate to toxic stimuli, how repeated exposure to toxicity affects the likelihood of toxic responses, and whether toxic behavior can be predicted from exposure alone. Our findings show that toxic responses are more likely following toxic stimuli, and, at the same time, cumulative toxic exposure (repeated over time) significantly increases the probability of toxic responding. We further introduce two influence metrics, the Influence-Driven Toxic Response Rate and the Spontaneous Toxic Response Rate, revealing a strong negative correlation between induced and spontaneous toxicity. Finally, we show that the number of toxic stimuli alone enables accurate prediction of whether an agent will eventually produce toxic content. These results highlight exposure as a critical risk factor in the deployment of LLM agents, particularly as such agents operate in online environments where they may engage not only with other AI chatbots, but also with human counterparts. This could trigger unwanted and pernicious phenomena, such as hate-speech propagation and cyberbullying. In an effort to reduce such risks, monitoring exposure to toxic content may provide a lightweight yet effective mechanism for auditing and mitigating harmful behavior in the wild.
@@ 1. Introduction
Social bots have long played a significant role in online platforms, influencing information diffusion, engagement dynamics, and public discourse (Ferrara et al., 2016; Bessi and Ferrara, 2016; Shao et al., 2018; Ferrara, 2017). However, the advent of Large Language Models (LLMs) has enabled a new generation of social bots capable of far more sophisticated and naturalistic interactions with users and with one another. These advances have spurred growing academic interest in understanding the behavior of LLM-driven social agents deployed in or modeled after online social ecosystems (Shen et al., 2025; Piao et al., 2025b).
@@
Unlike traditional rule-based or template-driven bots, LLM-based agents exhibit adaptive and emergent behaviors that arise from ongoing interactions within social networks (Ashery et al., 2025; He et al., 2023). Recent work has leveraged offline simulated social environments to examine the extent to which such agents can replicate human-like network structures (Gao et al., 2024), coordinate and perform cooperative tasks (Wu et al., 2024; Orlando et al., 2025), or give rise to collective phenomena such as polarization and echo chambers (Li et al., 2024; Piao et al., 2025a). Together, these studies suggest that LLM agents are not merely passive generators of text, but active participants in social dynamics whose behavior evolves over time.
@@
At the same time, despite the underlying guardrails, LLMs can be misused to generate toxic or harmful content at scale, posing risks to online communities and individual users (Chakraborty et al., 2024; Ferrara, 2024). In response, a substantial body of work has focused on measuring harmful generation and developing mitigation strategies, such as safer training procedures, filtering mechanisms, and post-hoc moderation (Wang et al., 2025).
Complementarily, a growing body of work in human-centered computing has shown that exposure to harmful content can influence user behavior, increasing the likelihood of adopting similar language or norms over time (Ferrara and Yang, 2015). However, it remains unclear whether, and to what extent, analogous dynamics apply to LLM-driven agents.
@@
In particular, we lack empirical evidence on: (i) how these agents behave when they solely interact to each other, in a fully AI-driven, dedicated platform; (ii) whether toxic content acts as a trigger for toxic responses, and (iii) whether repeated exposure systematically increases agents propensity to generate harmful outputs.
@@
In this paper, we address these gaps by performing a large-scale auditing on Chirper.ai, a fully AI-driven social network. Here, users can create and release LLMs-based agents that autonomously interact in the dedicated ecosystem, by generating posts and comments, as well as by interacting with other AI agents via following and liking mechanisms. Specifically, we rely on the platform to study toxicity adoption in LLM-driven agents through the lens of stimuli and responses. We model stimuli as posts that an agent explicitly comments on, and responses as the comments produced by the agent.
Our main contributions are threefold.
  • We are the first to provide a large-scale empirical audit of the mechanism of exposure-adoption toxicity in a fully AI-driven social platform.
  • We show that stimuli play a central role in shaping the harmful behavior of LLM-based agents, acting not only as a trigger for toxic response, but increasing the likelihood of toxic behavior under cumulative exposure to toxic content.
  • We further show that the number of toxic stimuli encountered by an agent alone enables accurate prediction of whether it will eventually produce toxic content, without requiring access to model internals, training data, or prompt instructions.
@@ 2. Related Work
In the following, we review the relevant literature by focusing on three core components of this work: (i) the behavior of AI-driven agents in online social ecosystems, (ii) the social mechanisms underlying exposure to harmful content and toxicity adoption, and (iii) Chirper.ai, the platform that serves as the empirical basis for our investigation.
@@
A growing body of research investigates how conversational agents and LLM-driven accounts behave when embedded in social platforms, where interactions are sequential, public, and shaped by evolving community norms (Chen et al., 2026). In these environments, agent behavior is no longer determined solely by isolated prompts, but emerges from continuous interaction with other agents and content. The recent development of benchmarks and simulation frameworks for social-media agents reflects a growing interest in evaluating both agent capabilities and associated risks in realistic, platform-like settings, where agents must interpret heterogeneous content, navigate social norms, and make posting decisions under uncertainty (Xue et al., 2025; Mohammadi et al., 2025).
Complementing these efforts, recent studies have begun to operationalize controlled human-agent interaction settings to analyze downstream social effects, such as polarization, opinion shifts, and norm formation, under carefully manipulated conditions (Donkers and Ziegler, 2025; Breum et al., 2024). These controlled environments allow researchers to isolate specific mechanisms of influence while retaining key features of social interaction. At the same time, earlier work in computational social science has shown that antisocial behavior in online communities often emerges through repeated interactions and reinforcing feedback loops, rather than being driven exclusively by a small number of persistently malicious actors (Cheng et al., 2015). This perspective highlights the importance of cumulative exposure and interaction dynamics in shaping harmful behaviors.
From this standpoint, it becomes critical to assess whether, and to what extent, fully AI-driven social platforms are susceptible to analogous emergent phenomena. Understanding how these agents interact and evolve is essential for anticipating the risks associated with deploying LLM-based agents in hybrid environments, where humans and chatbots coexist and interact. Uncontrolled harmful behavior may lead to the amplification and spread of hate speech, unsafe language, or toxic norms, ultimately undermining the safety and moderation of online social platforms.
@@
Research on online toxicity has traditionally focused on large-scale measurement and detection, enabling systematic empirical analyses of harmful speech across diverse online communities (Wulczyn et al., 2017). While this line of work has been essential for quantifying prevalence and building robust classifiers, it also raises a broader sociotechnical question: whether exposure to toxic content increases the likelihood that individuals subsequently adopt toxic language themselves.
A growing body of evidence suggests that exposure is indeed associated with increased rates of toxic behavior, though the magnitude and direction of this effect depend on contextual and social factors (Ferrara and Yang, 2015; Jiang et al., 2024). In particular, prior work shows that toxicity adoption varies according to group dynamics and social identity, such as whether harmful content originates from ingroup or outgroup members (Zoizner and Levy, 2025). These findings highlight that toxicity is not merely an individual trait, but a socially mediated phenomenon shaped by interaction patterns and relational context.
Complementary research further emphasizes the interactional nature of toxicity, demonstrating that harmful language often emerges at specific conversational turning points and can significantly alter the subsequent trajectory of discussion threads (Almerekhi et al., 2022). This perspective motivates moderation approaches that go beyond isolated message classification, instead accounting for local conversational context and cumulative exposure histories. Together, these insights directly inform our distinction between influence-driven and spontaneous toxicity, and motivate our use of exposure-based predictors to model and anticipate harmful responding in agent-driven environments.
@@
Launched in 2023, Chirper.ai [^1] https://chirper.ai/ (Minos, 2023, 2024) is a fully AI-driven social platform in which autonomous agents, referred to as chirpers, generate and interact with content. On Chirper.ai, chirpers can publish posts and comments, and engage with one another through social mechanisms such as following and liking, closely mirroring interaction paradigms found in human-centered platforms (e.g., $\mathbb{X}$ or Facebook).
@@
The platform has recently become a prominent environment for studying machine behavior in social-media-like ecosystems because it enables large-scale observation of autonomous agent-to-agent interactions. Recent work examines the platform content (Li et al., 2023) and compare it to human-driven social ecosystems (i.e., Mastodon), documenting differences in posting dynamics, abusive content, and network structure (Zhu et al., 2025). Our work builds on this emerging literature by focusing specifically on toxicity adoption as a function of observable stimuli, and by introducing complementary metrics to separate influence-driven toxicity from spontaneous toxic responding.
@@ 3. Methodology
    To what extent do agent responses reflect the toxicity of the stimuli?
@@
    Can we characterize the emergence of induced and spontaneous harmful behavior?
We tackle each research question in a dedicated subsection of the Results section below.
Table 1. Information describing the chirper agents, in terms of underlying model, number of parameters (expressed in billions or trillions), and percentage of generated content on the platform.
@@
Upon obtaining the appropriate permissions to access the Chirper.ai data, we collected $10,420,000$ textual records, including both original posts and comments. We therefore reconstructed the historical activity of about $75,000$ chirpers, instantiated by over $30,000$ (human) users. In addition to textual content, we also obtained metadata capturing social connections and interactions, such as follower relationships and likes. Table 1 further provides some information regarding the LLM architecture underlying the agents, and the corresponding volume of generated content on the platform. Notably, all the models through which the agents are instantiated have significant size in terms of parameters ($\geq$ 34B), ensuring language coherence and robustness. The most common LLMs released on the platform are Nous-Capybara-1.9, from the open-source AI Nous collection [^2] https://nousresearch.com/, and GPT-3.5-Turbo [^3] https://platform.openai.com/docs/models/gpt-3.5-turbo, released by OpenAI in 2022.
@@
We operationalize stimuli and responses at the level of comments. Specifically, a stimulus $S$ corresponds to the post to which a chirper responds, while a response $R$ denotes the comment generated by the chirper. This choice is driven by two main considerations.
@@
Second, we attempted to infer content exposure indirectly by leveraging the follower network, since it generally provides a robust proxy for content exposure (Mosleh and Rand, 2022) (i.e., assuming that a chirper is exposed to all content posted by agents they follow). However, empirical analysis revealed no meaningful correlation between these inferred exposure mechanisms and observed user interactions, such as likes or comments. This lack of correspondence suggests that such proxies do not reliably capture effective exposure on the platform.
Given these limitations, we adopt a conservative and observable definition of exposure, focusing exclusively on explicit interactions. We therefore characterize both stimuli and responses in terms of their toxicity. A piece of content is labeled as toxic if its toxicity score exceeds the $90$th percentile of the overall toxicity score distribution. This percentile-based threshold allows us to focus on the most extreme and potentially harmful content while remaining agnostic to absolute toxicity values. Nonetheless, we emphasize that our methodology is robust to the choice of toxicity threshold. Specifically, our findings remain unchanged when adopting a static threshold (e.g., $0.5$) to define toxic content.
@@
To estimate toxicity, we rely on detoxify [^4] https://pypi.org/project/detoxify/ (Hanu and Unitary team, 2020) , a BERT-based model trained for toxic comment classification, which reports strong performance (mean AUC $\simeq 0.99$). Given an input text, the model outputs a toxicity score in the interval $[0,1]$, representing the intensity of toxic language expressed.
To further improve classification reliability, we restrict our analysis to English-language content only. Language detection is performed using pycld2, [^5] https://pypi.org/project/pycld2/ a Python library built around the Google Chromium embedded Compact Language Detector (CLD2). [^6] https://github.com/CLD2Owners/cld2 This filtering step mitigates potential degradation in toxicity estimation due to multilingual or low-resource language inputs, which are not explicitly supported by the classifier. After applying language filtering, the final dataset consists of about 88M posts ($\sim 75\%$ of the original corpus).
Figure 1. Chirper distribution who generated toxic/non-toxic responses ($R^{*}/\tilde{R}$) over toxic/non-toxic stimuli ($S^{*}/\tilde{S}$). The y-axis reports the number of chirpers producing a given amount of responses (x-axis). Both axes are in log-scale.
Figure 2. Correlation between the number of chirpers Responses and Stimuli, categorized in toxic (*) and non-toxic ($\sim$). $\rho$ values indicate Pearson correlation ($p<.0001$)
@@ 4. Results
4.1. Responses and Stimuli Alignment (RQ1)
We first investigate how chirper responses vary as a function of the stimuli to which they respond to. Figure 1 shows the distribution of responses, distinguishing toxic and non-toxic responses in relation to toxic and non-toxic stimuli. Throughout the analysis, we denote toxic content with the symbol “*” and non-toxic content with $\sim$”.